What is the pH of a 0.180 M aqueous solution of potassium nitrite, KNO2 at 25 °C? (Ka for HNO2 = 4.5×10-4)
Answer #1

It is a salt of a weak acid and strong base, therefore it will undergo salt hydrolysis, anionic hydrolysis

pH of anionic hydrolysis is given by the foloofoll formula

pH = 1/2(pkw + pka + log(c))

Pkw = 14, pka = -logka = -log4.5×10-4 = 3.35

C = concentration = 0.18 M

pH = 1/2(14 + 3.35 + log(0.18))= 8.30

Ph = 8.30